Circular No. 8564

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

X-RAY FLASH IN M51 S. Immler, Goddard Space Flight Center, NASA, and USRA; and A.

Kong and W. H. G. Lewin, Massachusetts Institute of Technology, on behalf of a larger collaboration, report the detection of an x-ray flash in a 2136-s SWIFT XRT observation of M51 from July 6.231 UT (sequence 00030083011): “The significance of the detection is 6σ. Given an offset of the source from the tentative identification of the SN 2005cs progenitor (Richmond, IAUC 8555) of 8′′, and an XRT point-spread function of 18′′ (half-power diameter at 1.5 keV), this new source is probably unrelated to the supernova; however, based on the position alone, we cannot exclude the possibility that the emission is from SN 2005cs. The (backgroundsubtracted) count rate of the source is (7.4 ± 1.5) × 10^{−3} count/s, which corresponds to an unabsorbed 0.3–10-keV-band flux (averaged over the observation) of (3.3 ± 0.7) × 10^{−14} erg cm^{−2} s^{−1} and an average luminosity of (2.8 ± 0.6) × 10^{39} ergs/s, assuming a 10-keV thermal bremsstrahlung spectrum, an absorbing column density of 1.57 × 10^{20} cm^{−2} (Dickey and Lockman 1990, ARAA 28, 215), and a distance of 8.4 Mpc (Feldmeier et al. 1997, Ap.J. 479, 231). No x-ray source was detected 96 min later (sequence 00030083012, exposure 2031 s), as well as in all other SWIFT XRT observations of M51 on June 30, July 3, 5, 6, and 7, and in a previous 4800-s Chandra observation on 2003 Aug. 7. We encourage observations at other wavelengths.”

SUPERNOVA 2005ct IN NGC 207 A. V. Filippenko, F. J. D. Serduke, and S. Park, University of Cal-

ifornia, Berkeley; and D. C. Leonard, California Institute of Technology, report that inspection of CCD spectra (range 330–1000 nm), obtained on July 10 UT with the Shane 3-m telescope at Lick Observatory, shows that SN 2005ct (IAUC 8557) is a supernova, probably of type Ic (although type Ib cannot be ruled out at this time).

COMET P/2005 N3 (LARSON) As suspected on IAUC 8560, this comet is indeed of short period (P

= 6.7 yr); additional astrometry, elliptical orbital elements (T = 2005 Dec. 8.6 TT, e = 0.383, q = 2.199 AU, ω = 57^{o}.9, Ω = 298^{o}.5, i = 6^{o}.3, equinox 2000.0), and an ephemeris appear on MPEC 2005-N46.

2005 July 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green